A supermarket took a survey of 200 shoppers - 78 shoppers had hot-cross buns,92 shoppers had Easter eggs and 46 had neither. How many shoppers had a) both hot-cross buns and Easter eggs? b)had hot-cross buns but not Easter eggs?
Answer by ikleyn(52781)

.
A supermarket took a survey of 200 shoppers - 78 shoppers had hot-cross buns,92 shoppers had Easter eggs and 46 had neither.
How many shoppers had a) both hot-cross buns and Easter eggs? b)had hot-cross buns but not Easter eggs?
~~~~~~~~~~~~~~~~~~~~~~~

1.  200 - 46 = 154 was the number of those who had hot-cross buns OR had Easter eggs.



2.  78 + 92 - 154 = 16  was the number of those who had BOTH hot-cross buns AND Easter eggs.

    It is the answer to question a).


3.  78 - 16 = 62  was the number of those who had hot-cross buns but not Easter eggs.

    It is the answer to question b).

Solved.

All questions are answered.


For details and better understanding see the lesson
    - Counting elements in sub-sets of a given finite set
in this site.